4. Consider the boxplot shown below of the points earned on the 124 points of problems on the Fall 2018 STA2220 final exam for the 108 students in the 3 Tuesday/Thursday sections: -++Points o 10 20 30 40 50 60 70 80 90 100 110 120 130 (2 a) Describe the distribution as symmetric, left-skewed, or right-skewed. (2) b) About what proportion of the students scored at least half of the points possible on these problems? CS) c) Compute the values of the following statistics for these data. Show your computations. (Approximate the necessary values.) Range:__ Interquartile Range: Midrange:

asap please
Answer #1

(a) according to the box-plot given in the question, the distribution of the data is looked as a symmetric one because the distance between median and 1st quartile is almost same as the distance between median and 3rd quartile. So the data obtain in the given question coming from a symmetric distribution.

  (b) Total points is 124 so the half of the points = 62 around . and here median is nearly 70 so the proportion of student scored at least half of the points 0.55 approximately.

(c) The maximum point is nearly 125 and the minimum one is nearly 13 ..... so range = 125-13 = 112

1st quartile = 47 , 3rd quartile = 96, so the interquartile range = (3rd quartile - 1st quartile) = 96 - 47 = 49.

Mid range = ( 125+13 )/2 = arithmetic mean of maximum and minimum value of the data set = 69